A swimmer covers the 50-yard freestyle in a time of 0:2805 (28.05 sec), measured by electronic timer pads The double length of the pool measures as s . 50.0 ± 0.5 yards. The time can be taken as t 28.05 ±0.01 sec. The resulting speed, according to ves/t, Is then 1.78 yards/s Which of the following would be an acceptable way to give the uncertainty of the speed v? 0 Δu=0.5prd/s since the distance has the larger uncertainty. 。Aw = 0.5ー+ 50. Ord x 0.018 = 0.021/rd/8 28.05 (28.05 O Aw = 0.5yrd + 0.01 s = 0.51 yrd/s
Answer #1

v=\frac{s}{t} then proportional error in speed v is given by

\frac{\Delta v }{v}=\frac{\Delta s}{s}+\frac{\Delta t}{t}

hence uncertainty in speed is given by

\Delta v =\left ( \frac{\Delta s}{s}+\frac{\Delta t}{t} \right )v

Second option is more accurate way.

I will be glad to see your comment if you have any query and thumb up if you are satisfied. Thanks....